# Devils in Middle Layers of Large Vision-Language Models: Interpreting, Detecting and Mitigating Object Hallucinations via Attention Lens

### v2

@@ Abstract
Hallucinations in Large Vision-Language Models (LVLMs) significantly undermine their reliability, motivating researchers to explore the causes of hallucination. However, most studies primarily focus on the language aspect rather than the visual. In this paper, we address how LVLMs process visual information and whether this process causes hallucination. Firstly, we use the attention lens to identify the stages at which LVLMs handle visual data, discovering that the middle layers are crucial. Moreover, we find that these layers can be further divided into two stages: ”visual information enrichment” and ”semantic refinement” which respectively propagate visual data to object tokens and interpret it through text. By analyzing attention patterns during the visual information enrichment stage, we find that real tokens consistently receive higher attention weights than hallucinated ones, serving as a strong indicator of hallucination. Further examination of multi-head attention maps reveals that hallucination tokens often result from heads interacting with inconsistent objects. Based on these insights, we propose a simple inference-time method that adjusts visual attention by integrating information across various heads. Extensive experiments demonstrate that this approach effectively mitigates hallucinations in mainstream LVLMs without additional training costs.
@@ 1 Introduction
Building on the foundations of Large Language Models (LLMs), Large Vision-Language Models (LVLMs) [23, 24, 1, 4, 45, 38] have emerged as powerful tools for understanding visual inputs and generating textual responses aligned with user intent. Their multi-modal nature has enabled a large variety of vision-language tasks, e.g., image captioning, visual question answering, and visual dialog. Despite generating coherent textual responses, these models often exhibit undesirable object hallucinations – producing objects that do not exist in the visual input.
@@ 2 Related Work
Object Hallucination in LVLMs. Object hallucination [44, 42] is a prevalent and critical issue in current LVLMs where the model erroneously generates descriptions of non-existent objects in images. Previous studies to mitigate this issue have primarily focused on visual instruction fine-tuning [22, 17, 13, 40], incorporating external expert models [39, 5, 43, 37], and improving decoding strategies [16, 20, 25, 36, 8]. Despite numerous efforts, our understanding of the underlying mechanism driving these errors remains limited. Recent studies like OPERA [16] have pointed specific “anchor patterns” in text tokens that always align with the onset of hallucinated contents, and PAI [25] has identified “text inertia” as another contributing factor, as LVLMs reproduce identical hallucinated descriptions when solely input part of historical generated text. Contrary to these perspectives that focus on language bias, we investigate the causes of object hallucinations from the view of visual information processing through the lens of attention.
Interpretability in Foundation Models. In the field of NLP, numerous studies have analyzed LLMs to explore the internal model knowledge and understand model behavior in specific settings from the perspectives of attention maps [3, 41], neural activation patterns [15], and intermediate hidden states [31, 12]. The logit lens [26], which transforms the hidden states at each layer into the vocabulary space using the model’s own linear projector before softmax, offers insightful observations into the next token prediction process [10, 14] in LLMs. Our work adapts the logit lens method to analyze hidden states of image tokens, enabling us to interpret the LVLMs’ process of understanding the visual information via textual vocabulary. Few studies explore the internal mechanisms of LVLMs [2, 9], notable efforts include [27], which extends a unimodal causal tracing tool for studying neural mechanisms in BLIP’s image-conditioned text generation, and [28], which introduces a dictionary learning-based framework to extract multi-modal concepts for interpreting intermediate representations. Compared with these prior works, our study leverages the internal signals from attention weights to uncover the generation of hallucinated objects.
3 Understanding Object Hallucinations via Image Attention
This paper examines object hallucinations in LVLMs by analyzing attention weights and hidden states of image tokens. We start by introducing the generation process of LVLMs and the analytical tools, including visual attention ratio and logit lens. We then conduct case studies focusing on three key aspects: the processing of visual information, the attention patterns of object tokens, and the behavior of heads during object token generation.
@@ 3.1 Preliminary
Notation. The input to LVLMs is initialized with a sequence of $d$-dimensional image tokens $\{\mathbf{v}_{1},...,\mathbf{v}_{n}\}$ and instruction text (prompt) tokens $\{\mathbf{t}_{1},...,\mathbf{t}_{m}\}$. Typically structured as a Transformer decoder [35], the LVLM outputs responses in an autoregressive manner: at time step $k$, the model processes the initial input tokens $\{\mathbf{v}_{1},...,\mathbf{v}_{n},\mathbf{t}_{1},...,\mathbf{t}_{m}\}$, followed by the sequence of $k-1$ previously generated tokens $\{\mathbf{y}_{1},...,\mathbf{y}_{k-1}\}$, to predict the next token $\mathbf{y}_{k}$. Let $L$ denote the number of Transformer layers, each with $H$ heads. At layer $\ell\in[L]$, we denote the attention weights in head $h\in[H]$ as $\mathbf{A}_{k}^{(\ell,h)}\in\mathbb{R}^{a_{k}\times a_{k}}$, where $a_{k}=n+m+k-1$. Additionally, we define $\mathbf{v}^{\ell}_{i}$ and $\mathbf{y}^{\ell}_{k-1}$ as the hidden states for the image token $\mathbf{v}_{i}$ and generated token $\mathbf{y}_{k-1}$ at layer $\ell$, respectively. For convenience, we define the real and hallucinated object tokens as $\mathbf{y}_{\mathsf{r}}$ and $\mathbf{y}_{\mathsf{h}}$, respectively.
@@
where $\mathbf{A}_{k}^{(\ell,h)}(a_{k},i)$ represents the attention weights of the newly generated token $\mathbf{y}_{k}$ assigned to the image token $\mathbf{v}_{i}$. We use the VAR to quantify the extent of the token $\mathbf{y}_{k}$’s interaction with visual information: a higher VAR score indicating a greater contribution from image tokens during $\mathbf{y}_{k}$’s generation.
@@ 3.3 Finding 1: Middle Layers Matter for Visual Information Interaction
To assess the contribution of visual information to token generations, we compute the VAR score (Eq. 1) for real object tokens $y_{\mathsf{r}}$ across all layers $\ell\in[L]$ and heads $h\in[H]$. We calculate the mean of $\textrm{VAR}^{(\ell,h)}(y_{\mathsf{r}})$ for all real object tokens in our case dataset, with the results depicted in Fig. 2 (a). The visualization shows that the middle layers, i.e., 5-26 layers, continuously exhibit high attention weights to the image tokens, indicating that visual information interaction primarily occurs in the middle layers.
We further analyze how the model processes visual information by examining signals from the MHSA sublayers. Using the logit lens, we quantify the contribution of the MHSA sublayers to real token prediction as $\mathbf{p}_{\mathsf{o}}(\mathcal{V}|\mathbf{a}^{\ell}_{\mathsf{r}})$, where $\mathbf{a}^{\ell}_{\mathsf{r}}$ is the output of the MHSA sublayer at layer $\ell$, and $\mathsf{o}$ denotes the index of the real object token $y_{\mathsf{r}}$ in the vocabulary. Fig. 2 (b) reveals two patterns in the contributions from the MHSA sublayers: lower in 5-18 layers and higher in 19-26 layers, reflecting varied use of visual information.
@@ 3.4 Finding 2: Inactive Visual Attention in Middle Layers Implies Hallucination
Given that visual information is primarily processed in the middle layers, a natural question arises: does the model display unusual attention patterns in these layers when generating hallucinated object tokens? Fig. 4 presents the VAR distribution for real object tokens (‘bus’ and ‘car’) and hallucinated ones (‘people’ and ‘fire hydrant’) from the case of Fig. 3. We observe that the hallucinated object tokens demonstrate an inactive attention pattern during the visual information enrichment stage (5-18 layers) compared to the real ones. We conjecture that such pattern weakens the interaction between the object token and image tokens in this stage, limiting the propagation of visual information and potentially leading to hallucinations. To quantify this observation, we introduce a metric, Summed Visual Attention Ratio (SVAR), calculated by averaging VAR scores over all heads and summing across selected layers $[\ell_{s},\ell_{e}]$. Specifically, for a real object token $y_{\mathsf{r}}$ in $[\ell_{5},\ell_{18}]$, we compute:
$\textrm{SVAR}_{5\textrm{-}18}(y_{\mathsf{r}})\triangleq\frac{1}{H}\sum_{\ell=5}^{18}\sum_{h=1}^{H}\textrm{VAR}^{(\ell,h)}(y_{\mathsf{r}}).$ (3)
We conduct a statistical experiment on the $\textrm{SVAR}_{5\textrm{-}18}$ metric for both real and hallucinated tokens. The results, depicted in Fig. 5 (a), and the associated statistical test detailed in Sec. C.2, indicate a salient difference: the model assigns significantly higher attention weights to image tokens when generating real object tokens during visual information enrichment compared to hallucinated ones. A similar trend is observed in LLaVA-1.5-13B, presented in Fig. 5 (b).
@@ 3.4.1 Application: Object Hallucination Detection
Building upon our finding2, we apply the $\textrm{SVAR}_{5\textrm{-}18}$ metric to detect object hallucinations. We assess the efficacy of $\textrm{SVAR}_{5\textrm{-}18}$ in reflecting object hallucination using our case dataset, where a positive sample is a real object token and a negative sample is a hallucinated one. For comparison, we use the internal confidence metric recently proposed in [18] as a baseline. The internal confidence uses the maximum probability of the object token within all image hidden states $\{\mathbf{v}_{i}^{\ell}:i\in[n],\ell\in[L]\}$, mapped by the logit lens for detection. The qualitative results are presented in Fig. 6 and the results for LLaVA-1.5-13B are shown in Sec. C.1. Compared to internal confidence, using the simple $\textrm{SVAR}_{5\textrm{-}18}$ metric improves the AUROC by 8.82% and the AP by 3.53%, verifying the utility of our metric. The results further suggest that the middle layers provide crucial indicator information for object hallucination detection.
@@
Experimental setting. As shown in Fig. 8, we train a two-layer MLP by inputting the concatenated VAR scores across all heads in the layer range of interest $[\ell_{s},\ell_{e}]$ to classify the real and hallucinated object tokens. Specifically, the input vector $\mathbf{x}(y_{\mathsf{r}})$ for the real token $y_{\mathsf{r}}$ can be written as:
$\mathbf{x}(y_{\mathsf{r}})=[\textrm{VAR}^{(\ell_{s},1)},\textrm{VAR}^{(\ell_{s},2)},\dots,\textrm{VAR}^{(\ell_{e},H)}]^{\top}.$ (4)
@@ 4 Object Hallucination Mitigation
$\Delta$+3.3% -5.6% -21.8% -4.3% -38.3% -6.9% -7.8% -2.4% -19.5% -5.7%
Table 2: CHAIR hallucination evaluation and F1 results on three LVLMs with max new token set to 512. $\dagger$ denotes using the greedy decoding strategy. Our method outperforms other baselines and also preserves the richness of the descriptions.
@@ 4.2 Experimental Setup and Results
Results. Tab. 2 presents the experimental results for the selected LVLMs with max new tokens of 512. From the Tab. 2, our approach outperforms the three decoding strategies in reducing hallucinations and achieves an average reduction of 19.5% in $C_{S}$ and 5.7% in $C_{I}$, compared to the second best baseline. Compared to VCD, our method not only achieves superior hallucination mitigation performance but also better preserves the richness of the descriptions. Compared to PAI which requires the cost of once additional forward process for contrastive decoding, our approach achieves comparable or superior performance by only adjusting attention weights during inference, validating the efficacy of our method. Moreover, the consistent reduction of hallucinations across four different LVLMs confirms the generalizability of our findings. To evaluate the impact of generated token lengths, we conduct experiments on LLaVA-1.5-7B by varying the max new token in $\{64,128,256\}$. The results, reported in Tab. 3, show that our method consistently reduces hallucinations across different token lengths and achieves the best performance on average, demonstrating its robustness. We provide qualitative results in Sec. C.5.
@@ 5 Ablations and Discussions
Balance factor $\alpha$ sensitivity. To examine the influence of $\alpha$ os intervention efficacy, we vary its value in the range of $\{0.3,0.4,0.5,0.6,0.7\}$ and the results of CHAIR metrics and F1 score for the 7B version of both LLaVA-1.5 and MiniGPT-4 are shown in Fig. 9. We find that a lower $\alpha$ limits hallucination mitigation effectiveness, whereas a higher $\alpha$ compromises the richness of descriptions. These results indicate that an appropriate $\alpha$ value (e.g., 0.5 for both selected LVLMs) balances the reduction of hallucinations with the maintenance of detailed visual descriptions.

### v3

@@ Abstract
Hallucinations in Large Vision-Language Models (LVLMs) significantly undermine their reliability, motivating researchers to explore the causes of hallucination. However, most studies primarily focus on the language aspect rather than the visual. In this paper, we address how LVLMs process visual information and whether this process causes hallucination. Firstly, we use the attention lens to identify the stages at which LVLMs handle visual data, discovering that the middle layers are crucial. Moreover, we find that these layers can be further divided into two stages: “visual information enrichment” and “semantic refinement” which respectively propagate visual data to object tokens and interpret it through text. By analyzing attention patterns during the visual information enrichment stage, we find that real tokens consistently receive higher attention weights than hallucinated ones, serving as a strong indicator of hallucination. Further examination of multi-head attention maps reveals that hallucination tokens often result from heads interacting with inconsistent objects. Based on these insights, we propose a simple inference-time method that adjusts visual attention by integrating information across various heads. Extensive experiments demonstrate that this approach effectively mitigates hallucinations in mainstream LVLMs without additional training costs. [^1] Code: https://github.com/ZhangqiJiang07/middle_layers_indicating_hallucinations.
@@ 1 Introduction
Building on the foundations of Large Language Models (LLMs), Large Vision-Language Models (LVLMs) [27, 49, 4, 54, 44, 32, 51] have emerged as powerful tools for understanding visual inputs and generating textual responses aligned with user intent. Their multi-modal nature has enabled a large variety of vision-language tasks, e.g., image captioning, visual question answering, and visual dialog. Despite generating coherent textual responses, these models often exhibit undesirable object hallucinations – producing objects that do not exist in the visual input.
@@ 2 Related Work
Object Hallucination in LVLMs. Object hallucination [52, 48] is a prevalent and critical issue in current LVLMs where the model erroneously generates descriptions of non-existent objects in images, posing significant risks in high-stakes fields such as medical imaging [18, 23], sequential recommendation systems [15, 35], and autonomous driving [7]. Previous studies to mitigate this issue have primarily focused on visual instruction fine-tuning [26, 20, 14, 46, 53], incorporating external expert models [45, 5, 50, 43], and improving decoding strategies [19, 24, 28, 42, 9]. Despite numerous efforts, our understanding of the underlying mechanism driving these errors remains limited. Recent studies like OPERA [19] have pointed specific “anchor patterns” in text tokens that always align with the onset of hallucinated contents, and PAI [28] has identified “text inertia” as another contributing factor, as LVLMs reproduce identical hallucinated descriptions when solely input part of historical generated text. Contrary to these perspectives that focus on language bias, we investigate the causes of object hallucinations from the view of visual information processing through the lens of attention.
Interpretability in Foundation Models. In the field of NLP, numerous studies have analyzed LLMs to explore the internal model knowledge and understand model behavior in specific settings from the perspectives of attention maps [3, 47], neural activation patterns [17], and intermediate hidden states [36, 13]. The logit lens [29], which transforms the hidden states at each layer into the vocabulary space using the model’s own linear projector before softmax, offers insightful observations into the next token prediction process [11, 16] in LLMs. Our work adapts the logit lens method to analyze hidden states of image tokens, enabling us to interpret the LVLMs’ process of understanding the visual information via textual vocabulary. Few studies explore the internal mechanisms of LVLMs [1, 2, 10], notable efforts include [30], which extends an unimodal causal tracing tool for studying neural mechanisms in BLIP’s image-conditioned text generation, and [31], which introduces a dictionary learning-based framework to extract multi-modal concepts for interpreting intermediate representations. Compared to prior works, our study leverages the internal signals from attention weights to uncover the generation of hallucinated objects.
3 Understanding Object Hallucinations
In this section, we start by introducing the generation process of LVLMs and the analytical tools, including visual attention ratio and logit lens. We then conduct case studies focusing on three key aspects: the processing of visual information, the attention patterns of object tokens, and the behavior of heads during object token generation.
@@ 3.1 Preliminary
Notation. The input to LVLMs is initialized with a sequence of $d$-dimensional image tokens $\{\mathbf{v}_{1},...,\mathbf{v}_{n}\}$ and instruction text (prompt) tokens $\{\mathbf{t}_{1},...,\mathbf{t}_{m}\}$. Typically structured as a Transformer decoder [40], the LVLM outputs responses in an autoregressive manner: at time step $k$, the model processes the initial input tokens $\{\mathbf{v}_{1},...,\mathbf{v}_{n},\mathbf{t}_{1},...,\mathbf{t}_{m}\}$, followed by the sequence of $k{-}1$ previously generated tokens $\{\mathbf{y}_{1},...,\mathbf{y}_{k-1}\}$, to predict the next token $\mathbf{y}_{k}$. Let $L$ denote the number of Transformer layers, each with $H$ heads. At layer $\ell{\in}[L]$, we denote the attention weights in head $h{\in}[H]$ as $\mathbf{A}_{k}^{(\ell,h)}{\in}\mathbb{R}^{a_{k}\times a_{k}}$, where $a_{k}{=}n{+}m{+}k{-}1$. Additionally, we define $\mathbf{v}^{\ell}_{i}$ and $\mathbf{y}^{\ell}_{k-1}$ as the hidden states for the image token $\mathbf{v}_{i}$ and generated token $\mathbf{y}_{k-1}$ at layer $\ell$, respectively. The real and hallucinated object tokens are noted as $\mathbf{y}_{\mathsf{r}}$ and $\mathbf{y}_{\mathsf{h}}$, respectively.
@@
where $\mathbf{A}_{k}^{(\ell,h)}(a_{k},i)$ represents the attention weights of the newly generated token $\mathbf{y}_{k}$ assigned to the image token $\mathbf{v}_{i}$. VAR quantifies the extent of the token $\mathbf{y}_{k}$’s interaction with visual information: a higher VAR score indicating a greater contribution from image tokens during $\mathbf{y}_{k}$’s generation.
@@ 3.3 Finding 1: Middle Layers Matter for Visual Information Interaction
To assess the contribution of visual information to token generations, we compute the VAR score (Eq. 1) for real object tokens $\mathbf{y}_{\mathsf{r}}$ across all layers $\ell{\in}[L]$ and heads $h{\in}[H]$. We calculate the mean of $\textrm{VAR}^{(\ell,h)}(\mathbf{y}_{\mathsf{r}})$ for all real object tokens in our case dataset, with the results depicted in Fig. 2 (a). The visualization shows that the middle layers, i.e., 5-26 layers, continuously exhibit high attention weights to the image tokens, indicating that visual information interaction primarily occurs in the middle layers.
We further analyze how the model processes visual information by examining signals from the MHSA sublayers. Using the logit lens, we quantify the contribution of the MHSA sublayers to real token prediction as $\mathbf{p}_{\mathsf{o}}(\mathcal{V}|\mathbf{a}^{\ell}_{\mathsf{r}})$, where $\mathbf{a}^{\ell}_{\mathsf{r}}$ is the output of the MHSA sublayer at layer $\ell$, and $\mathsf{o}$ denotes the index of the real object token $\mathbf{y}_{\mathsf{r}}$ in the vocabulary. Fig. 2 (b) reveals two patterns in the contributions from the MHSA sublayers: lower in 5-18 layers and higher in 19-26 layers, reflecting varied use of visual information.
@@ 3.4 Finding 2: Inactive Visual Attention in Middle Layers Implies Hallucination
Given that visual information is primarily processed in the middle layers, a natural question arises: does the model display unusual attention patterns in these layers when generating hallucinated object tokens? Fig. 4 presents the VAR distribution for real object tokens (‘bus’ and ‘car’) and hallucinated ones (‘people’ and ‘fire hydrant’) from the case of Fig. 3. We observe that the hallucinated object tokens demonstrate an inactive attention pattern during the visual information enrichment stage (5-18 layers) compared to the real ones. We conjecture that such pattern weakens the interaction between the object token and image tokens in this stage, limiting the propagation of visual information and potentially leading to hallucinations. To quantify this observation, we introduce a metric, Summed Visual Attention Ratio (SVAR), calculated by averaging VAR scores over all heads and summing across selected layers $[\ell_{s},\ell_{e}]$. Specifically, for a real object token $\mathbf{y}_{\mathsf{r}}$ in $[\ell_{5},\ell_{18}]$, we compute:
$\textrm{SVAR}_{5\textrm{-}18}(\mathbf{y}_{\mathsf{r}})\triangleq\frac{1}{H}\sum_{\ell=5}^{18}\sum_{h=1}^{H}\textrm{VAR}^{(\ell,h)}(\mathbf{y}_{\mathsf{r}}).$ (3)
We conduct a statistical experiment on the $\textrm{SVAR}_{5\textrm{-}18}$ metric for both real and hallucinated tokens. The results, depicted in Fig. 5 (a), and the associated statistical test detailed in Sec. C.2, indicate a salient difference: the model assigns significantly higher attention weights to image tokens when generating real object tokens during visual information enrichment compared to hallucinated ones. Fig. 5 (b) shows a similar trend in LLaVA-1.5-13B.
@@ 3.4.1 Application: Object Hallucination Detection
Building upon our Finding 2, we apply the $\textrm{SVAR}_{5\textrm{-}18}$ metric to detect object hallucinations. We assess the efficacy of $\textrm{SVAR}_{5\textrm{-}18}$ in reflecting object hallucination using our case dataset, where a positive sample is a real object token and a negative sample is a hallucinated one. For comparison, we use the internal confidence metric recently proposed in [21] as a baseline. The internal confidence uses the maximum probability of the object token within all image hidden states $\{\mathbf{v}_{i}^{\ell}:i{\in}[n],\ell{\in}[L]\}$, mapped by the logit lens for detection. The qualitative results are presented in Fig. 6 and the results for LLaVA-1.5-13B are shown in Sec. C.1. Compared to internal confidence, using the simple $\textrm{SVAR}_{5\textrm{-}18}$ metric improves the AUROC by 8.82% and the AP by 3.53%, verifying the utility of our metric. This further suggests that the middle layers provide crucial indicator information for object hallucination.
@@
Experimental setting. We train a two-layer MLP by inputting the concatenated VAR scores across all heads in the layer range of interest $[\ell_{s},\ell_{e}]$ to classify the real and hallucinated object tokens. Specifically, the input vector $\mathbf{x}(\mathbf{y}_{\mathsf{r}})$ for the real token $\mathbf{y}_{\mathsf{r}}$ can be written as:
$\mathbf{x}(\mathbf{y}_{\mathsf{r}})=[\textrm{VAR}^{(\ell_{s},1)},\textrm{VAR}^{(\ell_{s},2)},\dots,\textrm{VAR}^{(\ell_{e},H)}]^{\top}.$ (4)
@@ 4 Object Hallucination Mitigation
$\Delta$$\uparrow$3.3$\downarrow$5.6$\downarrow$21.8$\downarrow$4.3$\downarrow$38.3$\downarrow$6.9$\downarrow$7.8$\downarrow$2.4$\downarrow$19.5$\downarrow$5.7%
Table 2: CHAIR hallucination evaluation and F1 results on three LVLMs with max new token set to 512. $\dagger$ denotes using the greedy decoding strategy. Our method outperforms other baselines and also preserves the richness of the descriptions. $\Delta$denotes the relative performance improvement with respect to the second-best method.
@@ 4.2 Experimental Setup and Results
Results. Tab. 2 presents the experimental results for the selected LVLMs with max new tokens of 512. From the Tab. 2, our approach outperforms the three decoding strategies in reducing hallucinations and achieves an average reduction of 19.5% in $C_{S}$ and 5.7% in $C_{I}$, compared to the second-best baseline. Compared to VCD, our method not only achieves superior hallucination mitigation performance but also better preserves the richness of the descriptions. Compared to PAI which requires the cost of once additional forward process for contrastive decoding, our approach achieves comparable or superior performance by only adjusting attention weights during inference, validating the efficacy of our method. Moreover, the consistent reduction of hallucinations across four different LVLMs confirms the generalizability of our findings. To evaluate the impact of generated token lengths, we conduct experiments on LLaVA-1.5-7B by varying the max new token in $\{64,128,256\}$. The results, reported in Tab. 3, show that our method consistently reduces hallucinations across different token lengths and achieves the best performance on average, demonstrating its robustness. We provide qualitative results in Sec. C.7.
@@ 5 Ablations and Discussions
Balance factor $\alpha$ sensitivity. To examine the influence of $\alpha$ on intervention efficacy, we vary its value in the range of $\{0.3,0.4,0.5,0.6,0.7\}$ and the results of CHAIR metrics and F1 score for the 7B version of both LLaVA-1.5 and MiniGPT-4 are shown in Fig. 8. We find that a lower $\alpha$ limits hallucination mitigation effectiveness, whereas a higher $\alpha$ compromises the richness of descriptions. These results indicate that an appropriate $\alpha$ value (e.g., 0.5 for both selected LVLMs) balances the reduction of hallucinations with the maintenance of detailed visual descriptions.
@@ 6 Conclusion and Insights
Acknowledgment
This work is supported by the National Science Foundation of China (62376281, 62206048), the Natural Science Foundation of Jiangsu Province (BK20220819), and the Fundamental Research Funds for the Central Universities (2242024k30035).